Order: Ampicillin 1,500 mg IV q8h for a child who weighs 33 lb. The safe dosage range is 300 mg/kg/day in equally divided doses as an intermittent infusion q8h.

1. Calculate the child's weight in kg.

2. Determine the safe dose (mg per dose rounded to tenths).

3. Is the order safe or not and what is your plan of action?

4. If it is safe, dilute the drug to a total of 25 mL in a Buretrol set (microdrip) and infuse at 25 gtt/min. How many minutes would it take to infuse the dose?
Answer #1

Order dose Ampicillin 1,500 mg IV q8h

The child weight is 33 lbs

The safe dose range is 300 mg/kg/day in equally divided doses as an intermittent infusion q8h.

1.

1 kg = 2.2 lbs

So, 33 lbs \div 2.2 lbs/kg = 15 kg

Answer - The child weight is 15 kg.

2.

The safe dose range is 300 mg/kg/day

So, 300 mg/kg/day \times 15 kg = 4500 mg/day

Answer - The safe dose range for this child is 4500 mg/day.

3.

Order dose 1500 mg IV q8h

So, the total dose for the day will be 1500 mg \times 3 times/day = 4500 mg/day

Answer - The dose is safe.

4.

After dilution total volume is 25 ml

we are using a micro drip set (60 gtt/ml)

Order rate of infusion is 25 gtt/min

So, 25 ml \times 60 gtt/ml \div 25 gtt/min = 60 min

Answer - To infuse the dose it will take 60 minute.